OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). x-2y=0 4x-3y=15

OpenStudy (anonymous):

x-2y=0 x=2y 4(2y)-3y=15 8y-3y=15 5y=15 y=3 now put back into other eqn x-3y=0 x-3(3)=0 x=9 ordered pair of 9,3

OpenStudy (apoorvk):

okay @Kittyrae15 seems like you have trouble understanding how to deal with a system linear equations, right?

OpenStudy (anonymous):

mhm.

OpenStudy (anonymous):

are you still struggling? whats the problem?